Jasper Fruits Corporation wholesales peaches and oranges. Barbara Jasper is working with the company’s accountant to prepare next year’s budget. Ms. Jasper estimates that sales will increase 3 percent for peaches and 8 percent for oranges. The current year’s sales revenue data follow:

First Quarter Second Quarter Third Quarter Fourth Quarter Total
Peaches $ 235,000 $ 255,000 $ 315,000 $ 255,000 $ 1,060,000
Oranges 418,000 468,000 588,000 398,000 1,872,000
Total $ 653,000 $ 723,000 $ 903,000 $ 653,000 $ 2,932,000

Based on the company’s past experience, cost of goods sold is usually 65 percent of sales revenue. Company policy is to keep 20 percent of the next period’s estimated cost of goods sold as the current period’s ending inventory. (Hint: Use the cost of goods sold for the first quarter to determine the beginning inventory for the first quarter.)

Required

  1. Prepare the company’s sales budget for the next year for each quarter by individual product.

  2. If the selling and administrative expenses are estimated to be $630,000, prepare the company’s budgeted annual income statement.

  3. Ms.Jasper estimates next year’s ending inventory will be $35,500 for peaches and $56,500 for oranges. Prepare the company’s inventory purchases budgets for the next year, showing quarterly figures by product.
Answer #1

Part 1

First Quarter

Second Quarter

Third Quarter

Fourth Quarter

Total

Peaches

242020

262650

324450

262650

1091770

Orange

451440

505440

635040

429840

2021760

Total

$693460

$768090

$959490

$692490

$3113530

First quarter

Peaches = 235000*1.03 = 242020

Organe = 418000*1.08 = 451440

Second quarter

Peaches = 255000*1.03 = 262650

Orange = 468000*1.08 = 505440

Third quarter

Peaches = 315000*1.03 = 324450

Organe = 588000*1.08 = 635040

Fourth quarter

Peaches = 255000*1.03 = 262650

Organe = 398000*1.08 = 429840

Part 2

Jasper Fruits Corporation

Budgeted Annual Income Statement

Sales Revenue

3113530

Cost of Goods Sold

2023795

Gross Profit

1089735

Selling & Administrative Expenses

630000

Net Income

$459735

3113530*65% = 2023795

Part 3

Jasper Fruits Corporation

Inventory Purchases Budget

First Quarter

Second Quarter

Third Quarter

Fourth Quarter

Total

Peaches

Sales

242020

262650

324450

262650

1091770

Cost of Goods Sold

157313

170723

210893

170723

709652

Desired ending Inventory (20% * next quarter’s cost of goods sold)

20%

20%

20%

20%

Ending Inventory

34145

42179

34145

35500

35500

Total Inventory Required

191458

212902

245038

206223

745152

Less: Beginning Inventory

31463

34145

42179

34145

31463

Budgeted Inventory Purchases

$159995

$178757

$202859

$172078

$713689

Orange

Sales

451440

505440

635040

429840

2021760

Cost of Goods Sold

293436

328536

412776

279396

1314144

Desired ending Inventory (20% * next quarter’s cost of goods sold)

20%

20%

20%

20%

Ending Inventory

65707

82555

55879

56500

56500

Total Inventory Required

359143

411091

468655

335896

1370644

Less: Beginning Inventory

58687

65707

82555

55879

58687

Budgeted Inventory Purchases

$300456

$345384

$386100

$280017

$1311957